Numerical Analysis I, Fall 2017 (http://www.math.nthu.edu.tw/˜ wangwc/)

Preparation guide for Quiz 01

The exam problems will be closely related to your homework problems. Make sure you understand all of them.

1. Review basic programming skills including ’for loop’, ’while loop’, etc in your favorite programming language.

2. Study the details about floating point binary expressions, such as how to turn a decimal 0.2 into a binary expression and vice versa.

3. Review IEEE single (double) precision arithmetics such as, ”Why does it take 32 (64) bits to store a single (double) precision floating point number?”, ”How is machine epsilon obtained?”, etc.

4. Study the derivation of the upper bound for relative error cause by chopping and rounding.

5. Understand how floating errors are amplified through multiplication, division and addition of two numbers of the same sign.

6. Understand the source of loss of significance caused by subtraction and how to avoid them. Study the examples in the textbook.